# Persistent Topological Features in Large Language Models

### v1

@@ Abstract
Understanding the decision-making processes of large language models (LLMs) is critical given their widespread applications. Towards this goal, describing the topological and geometrical properties of internal representations has recently provided valuable insights. For a more comprehensive characterization of these inherently complex spaces, we present a novel framework based on zigzag persistence, a method in topological data analysis (TDA) well-suited for describing data undergoing dynamic transformations across layers. Within this framework, we introduce persistence similarity, a new metric that quantifies the persistence and transformation of topological features such as $p$-cycles throughout the model layers. Unlike traditional similarity measures, our approach captures the entire evolutionary trajectory of these features, providing deeper insights into the internal workings of LLMs. As a practical application, we leverage persistence similarity to identify and prune redundant layers, demonstrating comparable performance to state-of-the-art methods across several benchmark datasets. Additionally, our analysis reveals consistent topological behaviors across various models and hyperparameter settings, suggesting a universal structure in LLM internal representations.
@@ 1 Introduction
More recently, geometric studies made progress by introducing concepts like intrinsic dimension to characterize the manifold of internal representations and its evolution across layers [10, 11, 12]. These methods have been successfully applied to transformer models in various works [13, 14, 15]. One notable achievement of this approach has been to show the emergence of semantic knowledge and abstraction phases in the middle layers of models, rather than at the final layers, as might be intuitively expected. However, these approaches provide only a static view of internal representations and suffer limitations in tracking their changes across layers.
@@
A natural framework to address these limitations and to offer a more comprehensive characterization of the geometry of internal representations of neural networks is Topological Data Analysis (TDA). TDA is a set of unsupervised techniques that offers robust methods to describe the shape and structure of complex datasets. It has seen exponential growth with applications in computational biology [16], cosmology [17, 18]], personalized medicine [19], time-dependent data analysis [20], and machine learning [21], just to name a few. One prominent tool within TDA is persistent homology, which tracks the birth and death of topological features across different scales, thereby capturing the multiscale behavior of a point cloud. Several studies have proposed persistent homology to investigate neural networks and their internal representations (e.g. [22], [23, 24]).
@@
However, in the context of TDA applications, it has not yet been recognized that the internal representations of LLMs can essentially be viewed as dynamic point clouds evolving in time (layers). As pre-trained LLMs process inputs, they transform these point clouds within the representation space layer by layer, capturing essential features and relationships throughout the model’s depth. Thus, it is natural to interpret these transformations as an evolving discrete dynamical system. TDA offers a specific framework for sequences of point clouds, known as zigzag persistence, built to characterize time-varying point clouds and temporal networks.
Following these ideas, we build a new framework based on zigzag persistence to analyse internal representations in LLMs by characterizing the birth and death of topological features within the model’s layers. Our zigzag algorithm achieves the following results:
  • Persistence Similarity: We propose a new metric to measure which topological features persist across the layers of an LLM. Different than other similarity measures, persistence similarity tracks the entire trajectory of transformations between two layers.
@@
  • Model Pruning: By identifying layers with high persistence similarity, we prune redundant layers without significantly degrading performance, finding comparable results to state-of-the-art methods.
  • Consistency Across Models and Hyperparameters: Our findings indicate that the behavior of persistent topological features and their similarities are consistent across different models, layers, and choices of hyperparameters of the framework. This consistency suggests a degree of universality in the topological structure of LLM representations.
In summary, our framework presents a novel perspective by combining two fundamental elements: firstly, it provides a fine-grained geometric analysis of the internal representations through TDA; secondly, the zigzag persistence framework tracks the trajectory of topological features across layers. Distinct from traditional methods that solely compare representations at individual layers, our approach captures their entire evolutionary path, providing a richer understanding of how these features evolve and contribute to the model’s decision-making processes.
@@ 2 Related Work
Geometry and Topology of Internal Representations.
The manifold hypothesis suggests that high-dimensional data often lies on a lower-dimensional manifold [25]. The estimation of this approximated manifold, known as intrinsic dimension, changes dynamically in deep networks, expanding and contracting in ways that impact performance [10], learnability [12], and the network’s ability to generate flexible abstract data representations used for downstream tasks [11], [13]. Intrinsic dimension and neighbor composition analysis of internal representations of causal and masked transformer models helped in the localization of semantic information, and to highlight differences between real and artificial data [13, 14, 15]. Another approach to study the internal representation is to use topological methods of TDA. Studies on Convolutional Neural Networks (CNN) used topological descriptors to explore the shape of activation functions [26] or their relations to performance [23]. [24] introduced persistent homology dimension as an estimator of the intrinsic dimension of internal representations in CNNs, while [27] proposed a measure of similarity based on topological descriptors to compare representations. Betti numbers have been observed to remain stable across different datasets for the same architectures and to decrease as depth increases [28].
@@
Layer Pruning by Similarity in Large Language Models. Among existing methods to reduce the size of neural networks, layer pruning has gained particular relevance in the context of LLMs. The first applications to BERT models [40, 41, 42, 43] inspired a long series of experiments employing similar techniques [44, 45, 46, 47, 48, 49]. Many of these efforts base their methodology on similarity measures of internal representations, which have conveniently been summarized in a recent review [50]. In this work, we consider [6], which uses angular similarity, and [7], which uses Block-Influence similarity, as a reference point for comparison.
@@ 3 Method
In this section, we introduce the zigzag persistence framework, which we use to analyze the internal representations of LLMs pre-trained with an autoregressive loss. These models typically receive an input sequence of $n$ tokens (often representing a sentence) embedded in a $d$-dimensional space. The input is transformed across the layers of the network without altering the embedding dimension. Due to the autoregressive nature of these models, the representation of the last token in a sequence captures information about the entire sequence and is the only token used for predicting the next. As a result, we choose to focus on the final token representation at each layer when analyzing input sequences. Thus, our data points in the representation space are vectors in a $d$-dimensional space, $\{\mathbf{x}_{i}(\ell_{j})\}\in\mathbb{R}^{d}$, for $i=1,...,N_{\rm sentences}$ and $j=1,...,N_{\rm layers}$, with each vector representing a sentence at a specific layer. Drawing from applications of topological data analysis, and particularly zigzag persistence in similar contexts, we interpret each layer as a snapshot of a complex, high-dimensional, time-varying system evolving through different stages (layers).
3.1 Topological Data Analysis and Persistent Homology
@@
Filtrations can be generalized to a more flexible structure called a zigzag filtration. Unlike a standard filtration, a zigzag filtration allows the sequence of complexes to move both forward and backward, meaning that inclusions between complexes can reverse at certain steps. We take this approach in our study to track the evolution of the internal representations across layers, rather than at a fixed snapshot, as done in traditional persistent homology implementations. In this sense, our parameter is not a distance/coarsening scale, but a discrete time scale represented by the layer number. We track topological features as they are formed and destroyed along the layers of the model and we statistically characterize these changes to describe a complex series of transformations in high-dimensional space. Differently than standard persistent homology, short- and long-lived features represent how the model dynamically evolves. Short-lived features indicate a high rate of rearrangement of the points $x_{i}$ between adjacent layers, while long-lived features suggest a phase of retention of (relative) positions in the model. This is a crucial point in our analysis, as it provides a novel tool to geometrically interpret the model’s internal representations. We now outline the main steps of the zigzag algorithm, leaving a rigorous mathematical formulation to Appendix A.
Figure 1: A schematic representation of the zigzag algorithm.
3.2 The zigzag algorithm
@@
We aim to study internal representations by tracking statistical changes in the formation of $p$-dimensional holes, or $p$-cycles, generated by connecting nearby data points within each layer $\ell_{i}$. As introduced above, the first ingredient for a TDA formulation is a criterion for connecting points of the dataset. In this regard, we construct a k-Nearest Neighbors graph $G_{\ell_{i}}=(V_{\ell_{i}},E_{\ell_{i}})$ at every layer $\ell_{i}$, where the number $k_{\rm NN}$ of neighbors is a fixed hyperparameter. To exploit the knowledge that the manifold on which the data lie is typically much smaller than the high-dimensional ambient space, we extend the dimension of the graph by filling higher-dimensional simplices. More precisely, we fill a simplex when its boundary, composed of lower-dimensional simplices (such as vertices and edges), is complete. In particular, we consider a triangle as filled when it has three vertices with pairwise connections. Similarly, a tetrahedron is filled when four vertices are all interconnected by edges, totaling six edges. This concept extends to higher dimensions up to a specified maximum dimension $m$. Thus, in each layer, we construct the simplicial complex $K_{\ell_{i}}$ defined by:
$K_{\ell_{i}}=\bigcup_{S\subseteq V_{\ell_{i}}}\big{\{}S\hskip 5.69054pt\big{|}\hskip 5.69054pt\forall x_{s},x_{l}\in S,\hskip 2.84526pt(x_{s},x_{l})\in E_{\ell_{i}}\hskip 2.84526pt\text{and}\hskip 2.84526pt|S|\leq m+1\big{\}}.$ (1)
@@
where we define $L\equiv N_{\rm layers}$ for conciseness. This sequence represents our zigzag filtration, denoted by $\Phi$. This filtration is the second ingredient needed to define persistent homology. We thus define a notion of birth and death of $p$-dimensional topological features, also denoted as $p$-cycles, with $p=0,...,m-1$, being $m$ the maximum dimension at which we expand the graph. Throughout this work, we choose $m=4$, which implies that the $p$-cycles are well defined up to dimension $p=3$. These cycles can be thought of as holes in their respective dimension. We can track the persistence of these cycles as they appear in a given layer when a group of points exhibits a particular proximity and distribution in the complex and disappear at a subsequent layer when some points have moved apart, causing the cycle to vanish. We illustrate the idea in Figure 1. The output of the zigzag algorithm is then a multiset of birth-death pairs $[\rm birth,\rm death]$ [^1] The repetition of a pair $[\rm birth,\rm death]$ indicates that multiple cycles in dimension $p$ have been created and destroyed in correspondence of the same layers., known as the persistence diagram

### v3

@@ Abstract
Understanding the decision-making processes of large language models is critical given their widespread applications. To achieve this, we aim to connect a formal mathematical framework—zigzag persistence from topological data analysis —with practical and easily applicable algorithms. Zigzag persistence is particularly effective for characterizing data as it dynamically transforms across model layers. Within this framework, we introduce topological descriptors that measure how topological features, $p$-dimensional holes, persist and evolve throughout the layers. Unlike methods that assess each layer individually and then aggregate the results, our approach directly tracks the full evolutionary path of these features. This offers a statistical perspective on how prompts are rearranged and their relative positions changed in the representation space, providing insights into the system’s operation as an integrated whole. To demonstrate the expressivity and applicability of our framework, we highlight how sensitive these descriptors are to different models and a variety of datasets. As a showcase application to a downstream task, we use zigzag persistence to establish a criterion for layer pruning, achieving results comparable to state-of-the-art methods while preserving the system-level perspective.
@@ 1 Introduction
More recently, geometric studies made progress by introducing concepts like intrinsic dimension to characterize the manifold of internal representations and its evolution across layers [10, 11, 12]. These methods have been successfully applied to transformer models in various works [13, 14, 15, 16, 17]. One notable achievement of this approach has been to show the emergence of semantic knowledge and abstraction phases in the middle layers of models, rather than at the final layers, as might be intuitively expected.
However, these approaches provide only a static view of internal representations and suffer limitations in tracking their changes across layers.
@@
A natural framework to address these limitations and to offer a more comprehensive characterization of the geometry of internal representations of neural networks is Topological Data Analysis (TDA). TDA is a set of unsupervised techniques that offers robust methods to describe the shape and structure of complex datasets. It has seen exponential growth with applications in computational biology [18], cosmology [19, 20]], personalized medicine [21], time-dependent data analysis [22], and machine learning [23], just to name a few. One prominent tool within TDA is persistent homology, which tracks the birth and death of topological features across different scales, thereby capturing the multiscale behavior of a point cloud. Several studies have proposed persistent homology to investigate neural networks and their internal representations (e.g. [24], [25, 26, 27]).
@@
However, in the context of TDA applications, it has not yet been recognized that the internal representations of neural networks can essentially be viewed as point clouds dynamically evolving in time (layers). In the particular case of LLMs, as pre-trained models process inputs, they transform these point clouds within the representation space layer by layer, capturing essential features and relationships throughout the model’s depth. Thus, it is natural to interpret these transformations as an evolving discrete dynamical system. To address this problem, we exploit a TDA tool developed to characterize time-varying point clouds and temporal networks, known as zigzag persistence.
Our approach achieves the following results:
  • ZigZag Framework for LLMs: We build a fast and scalable pipeline to characterize the birth and death of topological features across transformer models’ layers. As new contributions in the context of zigzag applications, we introduce the k-Nearest Neighbors-based filtration, and we interpret layers as time snapshots, tracking the trajectory of features across layers.
  • Identification of Phases of Prompt Processing: Using interpretable topological descriptors, we characterize the model’s dynamical processing of prompts in representation space across layers. We use this characterization to identify four distinct phases as prompts move in representation space across layers: an initial phase with rapid rearrangement of positions, a middle phase characterized by stable, long-lived relations among prompts, a transition phase where the model refines these relations, and a final phase of new rearrangements preparing data for output.
@@
  • Model Pruning: As a showcase downstream task, we use our descriptors to define a criterion to prune layers without significantly degrading performance, finding comparable results to state-of-the-art methods.
Our topological descriptors show quantitatively different results but qualitatively similar across models, datasets, and choices of zigzag hyperparameters. This proves their expressivity and simultaneously shows a degree of universality in the topological structure of LLM representations.
@@ 2 Related Work
Topology of Internal Representations. TDA has been extensively used in machine learning (see [23] for a recent review). In the context of studying internal representation, studies on Convolutional Neural Networks (CNN) used topological descriptors to explore the shape of activation functions [28] or their relations to performance [25]. [27] introduced persistent homology dimension as an estimator of the intrinsic dimension of internal representations in CNNs, while [29] proposed a measure of similarity based on topological descriptors to compare representations. Betti numbers have been observed to remain stable across different datasets for the same architectures and to decrease as depth increases [30].
@@
Distinguishing Transformer Stages. Recent works have studied the geometry of internal representations to identify distinct stages [^1] While previous work has frequently used the word “stage” in this context, we prefer “phase” to emphasize the continuous, evolving nature of the process. in the way pre-trained transformer models process input across layers. For instance, [13] demonstrated that transformer models exhibit a phase transition in the middle layers, characterized by a spike in intrinsic dimensionality, which correlates with the emergence of semantic and syntactic abstractions. Similarly, [15] highlighted that middle layers play a crucial role in compressing input representations into lower-dimensional manifolds, enabling the model to generalize and handle complex linguistic tasks. Subsequent work has confirmed and expanded these findings in different ways [42, 43, 44].
Layer Pruning in Large Language models. Among existing methods to reduce the size of neural networks, layer pruning has gained particular relevance in the context of LLMs. The first applications to BERT models [45, 46, 47, 48] inspired a long series of experiments employing similar techniques [49, 50, 51, 52, 53, 54]. Many of these efforts base their methodology on similarity measures of internal representations, which have conveniently been summarized in a recent review [55]. In this work, we consider [6], which uses angular similarity, and [7], which uses Block-Influence similarity, as a reference point for comparison.
@@ 3 Method
In this section, we introduce the zigzag persistence framework, which we use to analyze the internal representations of LLMs pre-trained with an autoregressive loss. These models typically receive an input sequence of $n$ tokens (often representing a sentence) embedded in a $d$-dimensional space. The input is transformed across the network layers without altering the embedding dimension. Due to the autoregressive nature of these models, the representation of the last token in a sequence captures information about the entire sequence and is used to predict the next. As a result, we choose to focus on the last token representation of each sequence at each layer. Thus, our point cloud is represented by last tokens embeddings, i.e. vectors of the form $\{\mathbf{x}_{i}(\ell_{j})\}\in\mathbb{R}^{d}$, for $i=1,...,N_{\rm sentences}$ and $j=1,...,N_{\rm layers}$. These last tokens are extracted from a variety of datasets and serve as an observational probe of how the model processes input.
3.1 Topological Data Analysis
@@
Filtrations can be generalized to a more flexible structure called a zigzag filtration. Unlike a standard filtration, a zigzag filtration allows the sequence of complexes to move both forward and backward, meaning that inclusions between complexes can reverse at certain steps. We take this approach in our study to track the evolution of the internal representations across layers, rather than at a fixed snapshot, as done in traditional persistent homology implementations. In this sense, our parameter is not a distance/coarsening scale, but a discrete time scale represented by the layer number. We track topological features as they are formed and destroyed along the model layers and statistically characterize these changes to describe a complex series of transformations in high-dimensional space. Differently than standard persistent homology, short- and long-lived features represent how the model dynamically evolves. Short-lived features indicate a high rate of rearrangement of the points between adjacent layers, while long-lived features suggest a phase of retention of (relative) positions across several layers. This is a crucial point in our analysis, as it provides a novel tool to interpret how the model processes different inputs by moving them and changing their relative positions in the representation space. Recent computational advances have allowed feasible implementation of these methods for analyzing time-varying point clouds within the broader machine learning community. For instance, efficient algorithms such as the fast zigzag persistence method introduced by [59] have enabled scalable analysis of evolving topological features, making persistence-based approaches much more accessible in large-scale applications. We now outline the main steps of the zigzag algorithm, leaving a rigorous mathematical formulation to Appendix A.
Figure 1: A schematic representation of the zigzag algorithm. It shows how zigzag can track the evolution of topological features over time, and the descriptors of this work are built upon them.
3.2 Zigzag Persistence for Layer Analysis
@@
We aim to study internal representations by tracking statistical changes in the formation of $p$-dimensional holes generated by connecting nearby data points within each layer $\ell_{i}$. As introduced above, the first ingredient for a TDA formulation is a criterion for connecting points. In this regard, we construct a k-Nearest Neighbors graph $G_{\ell_{i}}=(V_{\ell_{i}},E_{\ell_{i}})$ at every layer $\ell_{i}$, where the number $k_{\rm NN}$ of neighbors is a fixed hyperparameter (see [60] for a previous use of a $k_{\rm NN}$-based filtration). To explore higher-order relations among points, we extend the dimension of the graph by filling higher-dimensional simplices. More precisely, we fill a simplex when its boundary, composed of lower-dimensional simplices (such as vertices and edges), is complete. In particular, we consider a triangle as filled when it has three vertices with pairwise connections. Similarly, a tetrahedron is filled when four vertices are all interconnected by edges, totaling six edges. This concept extends to higher dimensions up to a specified maximum dimension $m$. Thus, in each layer, we construct the simplicial complex $K_{\ell_{i}}$ defined by:
$K_{\ell_{i}}=\bigcup_{S\subseteq V_{\ell_{i}}}\big{\{}S\hskip 5.69054pt\big{|}\hskip 5.69054pt\forall x_{s},x_{l}\in S,\hskip 2.84526pt(x_{s},x_{l})\in E_{\ell_{i}}$
$\hskip 2.84526pt\text{and}\hskip 2.84526pt|S|\leq m+1\big{\}}.$ (1)
@@
where we define $L\equiv N_{\rm layers}$ for conciseness. This sequence represents our zigzag filtration, denoted by $\Phi$. This filtration is the second ingredient needed to define persistent homology. We thus define a notion of birth and death of $p$-dimensional holes, with $p=0,...,m-1$, being $m$ the maximum dimension at which we expand the graph. Throughout this work, we choose $m=4$, which implies that the $p$-dimensional holes are well defined up to dimension $p=3$. We can track the persistence of these objects as they appear in a given layer when a group of points exhibits a particular proximity and distribution in the complex and disappear at a subsequent layer when some points have moved apart, causing them to vanish. We illustrate the idea in Figure 1.
Comparison to similar frameworks.
A distinguishing feature of our methodology is the choice of a $k_{\rm NN}$ filtration, whose stability was discussed in [60] in the context of persistent homology, though it was never applied to zigzag persistence. A notable effort in describing spatio-temporal networks similarly to this work is [40], where the main summary statistic (the rank invariant) involves calculating a 6-dimensional data vector (4 across layers and 2 across scale) and thus combines a variation of both a time and a scale parameter, using the Rips filtration. Varying also the scale parameter is worth investigating in this context, and the techniques in [40] would be a starting point for implementing it. The work [61] from the same authors is also related to our work since the maximal group diagram and the persistence clustergram (cfr. Figure 2) are “annotated” (with the representative topological features) barcodes. In this work, they fix a scale similar to our case.
Zigzag Persistence Diagram.
The output of the zigzag algorithm is then a multiset of birth-death pairs $[\rm b,\rm d]$ [^2] The repetition of a pair $[\rm b,\rm d]$ indicates that multiple holes in dimension $p$ have been created and destroyed in correspondence of the same layers., known as the persistence diagram